A 20 cm diameter jet of water strikes a curved vane with a velocity of 30 m/s. The inlet vane angle is zero and the outlet angle is 25°. Find the resultant force on the vane


(a) when the vane is fixed;


(b) when the vane is moving with a velocity of 15 m/s in the direction of jet.


1
Expert's answer
2021-08-26T01:05:06-0400

when the vane is fixed

"F_n = \rho A(V-u)^2sin\theta"

"F_n = 1000 \times \frac{\pi}{4}(0.2)^2\times 30 \times sin25 = 398.30 N"

when the vane is moving with a velocity of 15 m/s in the direction of jet.

"F_n = \rho A(V+u)^2sin\theta"

"F_n = 1000 \times \frac{\pi}{4}(0.2)^2\times 45 \times sin25 = 597.46 N"